Suppose an = 2n - 2*an-1 and a0=0

What is a2n?


1
Expert's answer
2021-10-25T16:09:32-0400

Write down a few first terms of a sequence:

"a_0=0\\
a_1=2-0=2\\
a_2=2^2-2\cdot2=0\\
a_3=2^3-0=2^3\\
a_4=2^4-2\cdot2^3=0"

it is easy to see that the even terms of a sequence are zero, so "a_{2n}=0"